Balerio Corporation's relevant range of activity is 8,000 units to 13,000 units. When it produces and sells 10,000 units, its average costs per unit are as follows: Average Cost per Unit Direct materials $ 7.20 Direct labor $ 3.80 Variable manufacturing overhead $ 2.10 Fixed manufacturing overhead $ 14.00 Fixed selling expense $ 2.60 Fixed administrative expense $ 1.60 Sales commissions $ 0.40 Variable administrative expense $ 0.30 Required: a. For financial reporting purposes, what is the total amount of product costs incurred to make 10,000 units? (Do not round intermediate calculations.) b. If 12,000 units are sold, what is the variable cost per unit sold? (Round "Per unit" answer to 2 decimal places.) c. If 12,000 units are sold, what is the total amount of variable costs related to the units sold? (Do not round intermediate calculations.) d. If the selling price is $18.80 per unit, what is the contribution margin per unit sold? (Round "Per unit" answer to 2 decimal places.) e. What incremental manufacturing cost will the company incur if it increases production from 10,000 to 10,001 units? (Round "Per unit"


a.Total product cost

b.Variable cost per unit sold

c.Total variable costs

d.Contribution margin per unit

e.Incremental manufacturing cost
Answer #1

a) Total product cost = (7.20+3.80+2.10+14)*10000 = 271000

b) Variable cost per unit sold = (7.20+3.80+2.10+0.40+0.30) = 13.80

c) Total variable cost = 13.80*12000 = 165600

d) Contribution margin per unit = 18.80-13.80 = 5

e) Incremental manufacturing cost = 7.20+3.80+2.10 = 13.10